OpenStudy (anonymous):

write the mixed number 2 1/4 as an improper fraction

OpenStudy (anonymous):

9/4

OpenStudy (anonymous):

\[2\frac{1}{4}\] to write it down as an improper fraction you'll have to multiply 4 with 2 and add 1 after that you divide by the denominator which is 4 ^_^: \[\frac{9}{4}\] do you understand it now? :)

OpenStudy (anonymous):

its (4x2) + 1

OpenStudy (anonymous):

yep ^_^

OpenStudy (anonymous):

Thats the numerator and the denominator remains 4

OpenStudy (anonymous):

do you get it now rsan? :)

OpenStudy (anonymous):

yea i did =] , i was writing the things down.

OpenStudy (anonymous):

lol, okay I have a question for you

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what is : \[3\frac{4}{5}\] as an improper fraction ^_^, no one answers but rsan please :)

OpenStudy (anonymous):

20/4

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

19/4 sorry =]

OpenStudy (anonymous):

think again :)read the notes

OpenStudy (anonymous):

what's your denominator? :)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

my dear :\[\frac{a}{b}\] a = numerator and b = denominator, so again what's your denominator?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

excellent, so what's the answer now ^_^?

OpenStudy (anonymous):

19/5

OpenStudy (anonymous):

bravo!, okay one more

OpenStudy (anonymous):

yay!!

OpenStudy (anonymous):

\[3 \frac{1}{4} + 2\frac{1}{2}\]=?

OpenStudy (anonymous):

don't rush :)

OpenStudy (anonymous):

hold on this might take a minute or two lol

OpenStudy (anonymous):

take your time ^_^

OpenStudy (anonymous):

how will i add 13/4 and 5/2? doesnt the denominator have to be the same?

OpenStudy (anonymous):

good, you're on the right track, so what will you do to get the denominator the same in this case ? :) do you have to multiply both or just one of the fractions? ^_^

OpenStudy (anonymous):

its 9/4

OpenStudy (anonymous):

no karrar lol, I gave her a question to solve ^_^

OpenStudy (anonymous):

HIM!!

OpenStudy (anonymous):

xD haha

OpenStudy (anonymous):

>_< sorry lol, please proceed, now add both factions ^_^

OpenStudy (anonymous):

fractions*

OpenStudy (anonymous):

you multiply just one freaction correct?

OpenStudy (anonymous):

yep and that is? :)

OpenStudy (anonymous):

5/2 multiply it and it becomes 10/4?

OpenStudy (anonymous):

excellent! now what's the answer?

OpenStudy (anonymous):

23/4

OpenStudy (anonymous):

braaaavoo!!! I would give you another medal, but I can only give one lol ^_^ now you understand the story, good luck!

OpenStudy (anonymous):

thank you again =]

OpenStudy (anonymous):

it's my pleasure :)